# Effects of Algorithmic Visibility on Conspiracy Communities: Reddit after Epstein’s ‘Suicide’

### v1

@@ 4.1. Behavioral and thematic responses (RQ1)
Figure 4. Shift in Empath category of core users’ posts spanning the arrest and the death of Jeffrey Epstein in r/conspiracy. Asterisks denote significance levels (***: $p<0.001$; **: $p<0.01$; *: $p<0.05$).
@@
Figure 5. Shift in Empath category between core and new users’ posts spanning the arrest and the death of Jeffrey Epstein in r/conspiracy. Asterisks denote significance levels (***: $p<0.001$; **: $p<0.01$; *: $p<0.05$).
@@
The lexical analysis with Empath reveals distinct thematic shifts both within the posts of core users over time and between core users and newcomers, as shown in Figure˜4 and Figure˜5. After Epstein’s death, core users increase their use of vocabulary related to crime and violence, while after the arrest, they reduce references to government, dispute, religion, and technology. These differences appear as statistically significant within-user shifts in the corresponding categories. The patterns show that established members change the topics and frames they emphasize around the two events, even when their overall posting volume remains high.
@@
Figure˜5 shows that, after Epstein’s death, new users use significantly more crime and violence terms and significantly fewer government terms than core users. After the arrest, the only significant shifts are negative: new users refer to government, politics, and religion less often than core users.
@@
Overall, the visibility event coincides with a sharp and lasting drop in how toxic core users’ posts are. This effect might be due to adaptation on the part of the core users to the influx of newcomers, whose toxicity level is reasonably lower, or due to the awareness of increased visibility due to having landed on Reddit’s homepage. At the same time, established members shift their language toward crime and violence and away from government- and religion-related themes, seemingly focusing more on the event at hand. Newcomers show similar topic choices but lean even more toward crime and less toward institutional politics than core users, which can be interpreted as reticence to engage with speculations and deeper conspiracy theories. These results show that the event changes both the tone and the main themes in the community’s discussions.
4.2. Community composition and user retention (RQ2)
@@
The second research question examines how homepage visibility affects users who join the community and how long they stay. A Kaplan-Meier survival analysis (Goel et al., 2010) reveals that the context of discovery is a strong predictor of retention. Users who join during the arrest period, which we assume corresponds to organic discovery, show higher survival probabilities at almost every point in time than users who join after the death event, corresponding to visibility-driven discovery.
@@
Figure 6. Kaplan-Meier survival curves for the four newcomer cohorts. From top to bottom over most of the time range, the curves are: arrest–engaged, arrest–not-engaged, death–engaged, and death–not-engaged. Solid and dashed line styles distinguish cohorts in addition to colour for accessibility.
@@
Figure˜6 shows that, among the arrest-era cohorts, users whose first post is about Epstein have the highest retention. Arrest-era users who start in non-Epstein threads leave more quickly but still outlast both death-era cohorts. Users who arrive via algorithmic visibility after the death event churn faster, regardless of whether their first post is about Epstein or about other conspiracy topics. The two death-era curves almost overlap, which indicates that discovery context matters more for retention than the specific initial topic.
Table 3. Integrated Brier Scores by cohort
@@
These findings reveal that users who discover the conspiracy community after Epstein’s arrest stay active much longer than those who arrive after his death, while post-death users churn quickly. This suggests that the context of discovery shapes not only how long people stay but also how diverse their engagement trajectories are. Users driven by algorithmic visibility are likely pushed by the exogenous stimulus of a collective attention spike; when the stimulus wanes, there is little endogenous engagement left with the community, so the users abandon it in a predictable way.
4.3. Linguistic integration patterns (RQ3)
The third research question asks whether the timing and context of entry influence how well new users integrate linguistically. The semantic distance analysis uses Sentence-BERT embeddings to track how close newcomer posts are to the weekly language of core users. Figure˜7 plots the weekly mean cosine distance for each cohort; values fall in a narrow band between roughly $0.66$ and $0.74$, so absolute effect sizes are modest, and the focus is on the relative ordering of cohorts over time.
Figure 7. Temporal evolution of semantic alignment for new user cohorts, measured as cosine distance from weekly core-user language centroids. Solid curves show three-week rolling means of weekly average distances; dots show unsmoothed weekly means. The light blue dashed curve indicates the arrest–not-engaged cohort, whose trajectory is sensitive to the choice of metric. The vertical dotted line marks Epstein’s death.
Users who arrive during the arrest period and discuss Epstein have the lowest cosine distance to core discourse and maintain this characteristic throughout the observation window. Their weekly trajectories show an initial convergence followed by stable alignment. Death-era cohorts, regardless of topic, stay at higher distances; they do not close the gap with the arrest–engaged cohort over the five months. The arrest–not-engaged cohort remains the farthest from the core and is the only cohort whose trajectory changes noticeably when using cross-entropy from bigram language models instead of cosine distance with SBERT (which is why it appears as a dashed line in Figure˜7).
Overall, the cosine distances vary within a narrow range, so the differences between cohorts are modest in absolute terms. Even so, users who arrive during the arrest period and post about Epstein use language that is consistently closest to that of core members. Death-era and non-engaged arrest-era users stay farther from core discourse over time, which suggests weaker linguistic integration.
5. Discussion
@@
We have measured toxicity over time and applied an Interrupted Time Series model to quantify changes among core users around Epstein’s death; this analysis revealed a sharp and sustained drop in toxicity immediately following the visibility spike. We have examined thematic dynamics using Empath lexical categories and found shifts toward crime- and violence-related language, alongside declines in government and religion framing. We have modeled newcomer retention through Kaplan-Meier survival curves and observed that users drawn by homepage exposure churn faster and more uniformly than those joining organically during the arrest period. Finally, we have evaluated linguistic integration using Sentence-BERT embeddings and bigram cross-entropy, showing that organic arrest-period newcomers converge most closely with core discourse, while post-death cohorts remain semantically distant throughout the study window.
@@
Our findings reveal several insights into how the algorithmic visibility event of being featured on Reddit’s homepage reshaped r/conspiracy. First, the toxicity analysis shows that core users reduced offensive language immediately after the spike and maintained lower toxicity levels without a clear rebound. This suggests that established members adjust their style when the community becomes more publicly exposed—whether through self-moderation, moderation enforcement, language mirroring, or social pressure from a broader audience.
@@
Second, the retention and IBS results indicate that discovery context functions as a compositional filter. Homepage visibility generates large influxes of short-lived participants who behave in a relatively uniform way, whereas organic discovery during the arrest period brought in fewer but more varied users, including those who remain active long-term. Combined with semantic distance findings, this result implies that the visibility spike expands the audience but not the pool of deeply integrated community members.
@@
Third, the semantic integration analysis shows stable and persistent differences between newcomers. Arrest–engaged users remain most aligned with core discourse, while death-era cohorts stay consistently farther away. Cross-entropy trajectories reinforce this pattern for three of the four cohorts. In this context, algorithmic visibility has an impact on who arrives and how they write, but it does not fundamentally shift the linguistic center of the community.
@@
These patterns contribute to debates on algorithmic exposure, conspiracy consumption, and radicalization—within the limits of observable posting activity. During this five-month window, most users drawn by homepage visibility do not remain active or converge toward core linguistic norms. We therefore find no evidence that incidental exposure alone generates large numbers of durable, linguistically integrated participants. Rather, users who organically discover the conspiracy community, and who might have an already-existing intrinsic motivation, are the ones that persist the longest (Corso et al., 2025b). This result does not preclude the existence of other forms of radicalization that occur privately, later, or off-platform; it simply indicates that such processes are not reflected in posting behavior in this dataset.
5.1. Limitations and future work
@@
Like any empirical work, this study is not exempt from limitations. The primary limitation is the observation window. The analysis covers only five months around the events related to Epstein and therefore focuses on a period of acute crisis rather than on longer arcs of community evolution. Integration and retention patterns may look different when the community is not under intense public attention, when other news dominates the public agenda, or when the subreddit experiences slower, more endogenous change. However, this event gives us an ideal natural experiment for the effects of algorithmic visibility that is invaluable to study the dynamics of opinion formation on social media.
@@
The availability and choice of specific measurement tools introduce further constraints. Toxicity and lexical categories come from models trained on general English and may miss context-specific slang, sarcasm, or coded language common in conspiracy communities. Sentence embeddings rely on a pre-trained SBERT model that may underweight very short posts, strip URLs and usernames into generic tokens, and compress different senses of the same word into a single vector. As a result, the semantic distance measures capture broad alignment rather than fine-grained ideological nuance. Using a mix of fine-tuned embeddings, large language models, and human validation could be a viable strategy to improve the reliability and accuracy of the current results.
@@
5.2. Ethical use of data and consent
This study analyzes discussion data from the subreddit r/conspiracy that was collected via the public Pushshift Reddit archives. The dataset consists solely of posts and comments that were already publicly visible on Reddit at the time of collection, and no direct contact with Reddit users took place. To protect user privacy, we conduct all analyses on de-identified records and report only aggregate results. No usernames, direct identifiers, or verbatim quotations that could enable re-identification are included in this paper. Rediscovery or amplification of harmful conspiratorial content is a potential risk in this domain; accordingly, we avoid reproducing sensitive text, refrain from linking specific posts or users, and restrict analysis to behavioral and linguistic aggregates. Data handling followed institutional ethical standards for research on public online communities.

### v2

@@ Results
Table 2: Interrupted time series regression estimates for core users’ toxicity.
@@
Figure 3: Shift in Empath category of core users’ posts spanning the arrest and the death of Jeffrey Epstein in r/conspiracy. Asterisks denote significance levels (***: $p<0.001$; **: $p<0.01$; *: $p<0.05$).
@@
Figure 4: Shift in Empath category between core and new users’ posts spanning the arrest and the death of Jeffrey Epstein in r/conspiracy. Asterisks denote significance levels (***: $p<0.001$; **: $p<0.01$; *: $p<0.05$).
@@
The lexical analysis with Empath reveals distinct thematic shifts both within the posts of core users over time and between core users and newcomers, as shown in Figure˜3 and Figure˜4. After Epstein’s death, core users increase their use of vocabulary related to crime and violence, while after the arrest, they reduce references to government, dispute, religion, and technology. These differences appear as statistically significant within-user shifts in the corresponding categories. The patterns show that established members change the topics and frames they emphasize around the two events, even when their overall posting volume remains high.
@@
Figure˜4 shows that, after Epstein’s death, new users use significantly more crime and violence terms and significantly fewer government terms than core users. After the arrest, the only significant shifts are negative: new users refer to government, politics, and religion less often than core users.
@@
Overall, the visibility event coincides with a sharp and lasting drop in how toxic core users’ posts are. This result might be due to several factors such as the adaptation on the part of the core users to the influx of newcomers, [^4] https://reddit.com/r/conspiracy/comments/cpb1ay whose toxicity level is reasonably lower, or due to the awareness of increased visibility due to having landed on Reddit’s homepage. At the same time, established members shift their language toward crime and violence and away from government- and religion-related themes, seemingly focusing more on the event at hand. Newcomers show similar topic choices but lean even more toward crime and less toward institutional politics than core users, which can be interpreted as reticence to engage with speculations and deeper conspiracy theories. These results show that the event changes both the tone and the main themes in the community’s discussions.
Community composition and user retention (RQ2)
@@
The second research question examines how homepage visibility affects users who join the community and how long they stay. A Kaplan-Meier survival analysis (Goel et al. 2010) reveals that the context of discovery is a strong predictor of retention. Users who join during the arrest period, which we use as a proxy for different a more organic discovery context, show higher survival probabilities at almost every point in time than users who join after the death event, corresponding to visibility-driven discovery.
@@
Figure 5: Kaplan-Meier survival curves for the four newcomer cohorts. From top to bottom over most of the time range, the curves are: arrest–engaged, arrest–not-engaged, death–engaged, and death–not-engaged. Solid and dashed line styles distinguish cohorts in addition to colour for accessibility.
@@
Figure˜5 shows that, among the arrest-era cohorts, users whose first post is about Epstein have the highest retention. Arrest-era users who start in non-Epstein threads leave more quickly but still outlast both death-era cohorts. Users who arrive during the amplified mainstream visibility after the death event churn faster, regardless of whether their first post is about Epstein or about other conspiracy topics. The two death-era curves almost overlap, which indicates that discovery context matters more for retention than the specific initial topic.
@@
These findings reveal that users who discover the conspiracy community after Epstein’s arrest stay active much longer than those who arrive after his death, while post-death users churn quickly. This suggests that the context of discovery shapes not only how long people stay but also how diverse their engagement trajectories are. Users whose activity coincides with sudden mainstream visibility are likely pushed by the exogenous stimulus of a collective attention spike; when the stimulus wanes, there is little endogenous engagement left with the community, so the users abandon it in a predictable way.
Linguistic integration patterns (RQ3)
The third research question asks whether the timing and context of entry influence how well new users integrate linguistically. The semantic distance analysis uses Sentence-BERT embeddings to track how close newcomer posts are to the weekly language of core users. Figure˜6 plots the weekly mean cosine distance for each cohort; values fall in a narrow band between roughly $0.66$ and $0.74$, so absolute effect sizes are modest, and the focus is on the relative ordering of cohorts over time.
Figure 6: Temporal evolution of semantic alignment for new user cohorts, measured as cosine distance from weekly core-user language centroids. Solid curves show three-week rolling means of weekly average distances; dots show unsmoothed weekly means. The light blue dashed curve indicates the arrest–not-engaged cohort, whose trajectory is sensitive to the choice of metric. The vertical dotted line marks Epstein’s death.
Users who arrive during the arrest period and discuss Epstein have the lowest cosine distance to core discourse and maintain this characteristic throughout the observation window. Their weekly trajectories show an initial convergence followed by stable and consistent lower semantic distance over time. Death-era cohorts, regardless of topic, stay at higher distances; they do not close the gap with the arrest–engaged cohort over the five months. The arrest–not-engaged cohort remains the farthest from the core and is the only cohort whose trajectory changes noticeably when using cross-entropy from bigram language models instead of cosine distance with SBERT (which is why it appears as a dashed line in Figure˜6).
Overall, the cosine distances vary within a narrow range, so the differences between cohorts are modest in absolute terms. Even so, users who arrive during the arrest period and post about Epstein use language that is consistently closest to that of core members. Death-era and non-engaged arrest-era users stay farther from core discourse over time, which suggests a possible weaker linguistic integration of this cohort. We observe that users with higher semantic distance also tend to have shorter retention, although causality cannot be established.
Discussion
@@
We have measured toxicity over time and applied an Interrupted Time Series model to quantify changes among core users around Epstein’s death; this analysis revealed a sharp and sustained drop in toxicity immediately following the visibility spike. We have examined thematic dynamics using Empath lexical categories and found shifts toward crime- and violence-related language, alongside declines in government and religion framing. We have modeled newcomer retention through Kaplan-Meier survival curves and observed that users who activate during homepage exposure churn faster and more uniformly than those joining organically during the arrest period. Finally, we have evaluated linguistic integration using Sentence-BERT embeddings and bigram cross-entropy, showing that organic arrest-period newcomers converge most closely with core discourse, while post-death cohorts remain semantically distant throughout the study window.
@@
Our findings reveal several insights into how the mainstream visibility event of being featured on Reddit’s homepage reshaped r/conspiracy. First, the toxicity analysis shows that core users reduced offensive language immediately after the spike and maintained lower toxicity levels without a clear rebound. This suggests that established members could have adjusted their style when the community becomes more publicly exposed—whether through self-moderation, moderation enforcement, language mirroring, or social pressure from a broader audience.
@@
Second, the retention and IBS results indicate that discovery context functions as a compositional filter. Homepage visibility is associated with large influxes of short-lived participants who behave in a relatively uniform way, whereas organic discovery during the arrest period brought in fewer but more varied users, including those who remain active long-term. Combined with semantic distance findings, this result suggests that the visibility spike expands the audience but not the pool of deeply integrated community members.
@@
Third, the semantic integration analysis shows stable and persistent differences between newcomers. Arrest–engaged users remain most aligned with core discourse, while death-era cohorts stay consistently farther away. Cross-entropy trajectories reinforce this pattern for three of the four cohorts. In this context, mainstream visibility has an impact on who arrives and how they write, but it does not fundamentally shift the linguistic center of the community.
@@
These patterns contribute to debates on mainstream exposure, conspiracy consumption, and radicalization—within the limits of observable posting activity. During this five-month window, most users who joined during homepage visibility do not remain active or converge toward core linguistic norms. We do not observe evidence that incidental exposure alone is associated with large numbers of durable, linguistically integrated participants. Rather, users who discover the conspiracy community during the arrest era, and who might have an already-existing intrinsic motivation, are the ones that persist the longest (Corso et al. 2025c). This result does not preclude the existence of other forms of radicalization that occur privately, later, or off-platform; it simply indicates that such processes are not reflected in posting behavior in this dataset.
Limitations and future work
@@
Like any empirical work, this study is not exempt from limitations. The primary limitation is the observation window. The analysis covers only five months around the events related to Epstein and therefore focuses on a period of acute crisis rather than on longer arcs of community evolution. Integration and retention patterns may look different when the community is not under intense public attention, when other news dominates the public agenda, or when the subreddit experiences slower, more endogenous change. However, this event gives us an ideal case study experiment for the effects of mainstream visibility that is invaluable to study the dynamics of opinion formation on social media.
@@
The availability and choice of specific measurement tools introduce further constraints. Toxicity and lexical categories come from models trained on general English and may miss context-specific slang, sarcasm, or coded language common in conspiracy communities. Sentence embeddings rely on a pre-trained SBERT model that may underweight very short posts, strip URLs and usernames into generic tokens, and compress different senses of the same word into a single vector. As a result, the semantic distance measures capture broad alignment rather than fine-grained ideological nuance. Using a mix of fine-tuned embeddings, large language models, and human validation could be a viable strategy to improve the reliability and accuracy of the current results. Moreover, we cannot distinguish whether these differences reflect post-entry linguistic adaptation or pre-existing similarity between newcomers and core users (selection effects).
@@
Ethical use of data and consent
This study analyzes discussion data from the subreddit r/conspiracy that was collected via the public Pushshift Reddit archives. The dataset consists solely of posts and comments that were already publicly visible on Reddit at the time of collection, and no direct contact with Reddit users took place. To protect user privacy, we conduct all analyses on de-identified records and report only aggregate results. No usernames, direct identifiers, or verbatim quotations that could enable re-identification are included in this paper. Rediscovery or amplification of harmful conspiratorial content is a potential risk in this domain; accordingly, we avoid reproducing sensitive text, refrain from linking specific posts or users, and restrict analysis to behavioral and linguistic aggregates. Data handling followed institutional ethical standards for research on public online communities. We leave the repository with the data and the code to replicate our analyses as follows: https://github.com/aattanasio/Paper.git
@@ Appendix A Appendix
Additional Figures
In Figure˜7 and Figure˜8 we show two different measurement of the exogenous event that caused the shock in the community: the death of Jeffrey Epstein. With the former, we compare the prevalence of the keywords ’Jeffrey Epstein’ and ’Reddit’ on Google Trends during the two periods of interest. We see how the death event provoked a much higher volume of researches. During the arrest era, both Jeffrey Epstein and Reddit see spikes of interest, possibly linked to users actively searching for further information or discussion on this topic on Reddit. While for the case of the death era, we see a significant spike of interest related to Jeffrey Epstein, but the same cannot be said for Reddit, which search volume remains mostly unchanged during that period.
Figure 7: Longitudinal Google Search prevalence of keywords: ‘Jeffrey Epstein’ and ‘Reddit’ during the arrest and death periods.
In Figure˜8 we show the unique active users during the period of our study. It’s evident how the death event brought a large volume of unique users to become active in the community, which is a different behavior from what can be observed a month prior for the arrest.
Figure 8: Unique users active on r/conspiracy during the period of our study.